# Gas contamination and mitigation in a 100 m3 / 10 bar argon TPC with optical readout: a viability study

### v1

@@ 4.3 Water contamination
For Teflon, outgassing rates of water are far too high compared to O2, so that if outgassing would not decrease over time then the necessary purity levels may never be achieved. Instead of assuming a constant value, we consider in this case the actual outgassing law based on eq. 7, with outgassing rates after 10 h taken from table 1: $3.5\times 10^{-8}$ $\frac{\textnormal{mbar}\cdot{\textnormal{l}}}{\textnormal{cm}^{2}\cdot{\textnormal{s}}}$ (‘optimistic’) and $10\times 10^{-8}$ $\frac{\textnormal{mbar}\cdot{\textnormal{l}}}{\textnormal{cm}^{2}\cdot{\textnormal{s}}}$ (‘typical’).
We concentrate on the case of typical outgassing and flow rates, and five different Teflon thicknesses, including the case where Teflon would be so thick so as to preserve a $t^{-1/2}$ behaviour all the time throughout. Based on the analysis in the previous section, the configuration with triple-holes has been chosen. Results are shown in Fig. 14. For thicknesses below 10 mm, a transition between the $t^{-1/2}$ and exponential regimes can be observed. As the outgassing rates of H2O are about 10 times larger than those of O2 in case of Teflon, the time scales involved are larger too. The horizontal line shows the 1 ppm-level, that is a plausible requirement to keep space resolution undistorted below 1 mm, based on the analysis presented in section 2.1. For a thin-Teflon reflector down to 0.5 mm thickness, coupled to an ESR reflective coating (see e.g. [52]), that land-mark value is achieved in 46 h, while it reaches up to 99 h as the Teflon becomes thicker. The result is sobering: when considering the displacement of air during filling and the time needed for water (oxygen) to fall below the concentration that makes its contribution negligible for tracking (attachment), about 6-10 days time would be typically needed. Well-defined cleaning protocols and surface treatment, as discussed for example in [37], together with an increase of the gas flow above the levels discussed here, or designing the system in compliance with the requirements needed for pumping, could plausibly reduce these times to the scale of about 1-2 days. On the other hand, additional (uncontrolled) outgassing sources would render the operation of the detector problematic.
@@
Additional studies are presented in Fig. 15 in order to illustrate the main dependencies. On the left-figure, the behaviour for different gases is shown for a fixed flow rate of 6.5 m3/h (argon: continuous lines; xenon: dashed lines). It can be seen how, despite the diffusion coefficient of H2O being slightly larger in argon by about 50% (table 1), the impact on the time-evolution of the impurities is almost indiscernible. When reducing the pressure down to 1 bar, the ×$\times 10$ increase in the diffusion coefficient accelerates the mixing as expected, however the reduction in the position of the maximum is a mere 30%. As the number of normal liters is accordingly reduced by a factor ×$\times 10$ in those conditions, the concentration of impurities increases in inverse proportion (the Ar and Xe lines at 1 bar have been downscaled by ×$\times 10$).
Fig. 15-right shows a comparison with the 1-D solutions for the case of time-dependent outgassing (see appendix). As for the case of constant outgassing, the results compare well with the 1-D model, despite the agreement is qualitatively worse. Artificially increasing the gas diffusion accelerates the mixing, but very mildly in the conditions studied (as indirectly shown in Fig. 15-left). The conclusion is once again that the velocity field in combination with gas diffusion can distribute the outgassed species uniformly over the chamber for the optimized ‘triplet’ distributor, as seen in Fig. 9. Also in this case, simulations for lower (‘optimistic’) outgassing rates display an exact down-scaling compared to the solutions presented in Fig. 15, as expected from the 1-D solutions.
@@ 4.4 Nitrogen contamination
Nitrogen can be removed from the system by flushing the main gas through ‘hot’ (temperature-activated) purifiers/getters, a common asset for the operation of noble gas detectors. The process of nitrogen purification becomes challenging, however, in case of operation under large gas flows (that could modify the getter temperature) and/or in the presence of molecular additives such as CF4 (that may react with the getter materials). Following conventional wisdom, and as long as a purification system compatible with Ar/CF4 operation is not demonstrated at the necessary scale, we will assume conservatively that N2 accumulates in the TPC as time passes.
@@
Fortunately, the tolerance to N2 is by far the largest of common outgassing species, up to 10 ppms (table 1). Based on the the time profile of the outgassing rate (eq. 7), the range of outgassing values in table 1, and the chamber volume, the evolution of N2 inside the TPC can be approximated as (see appendix):
$f_{\text{N}_{2}}\sim(3-9)\textnormal{ppms}\cdot\sqrt{t[\textnormal{days}]}.$ (26)
@@
with the pre-factor covering the range dubbed earlier as ‘optimistic’ (3 ppm) and ‘typical’ (9 ppm). In practice, this expression indicates that N2 can approach harmful concentrations pretty quickly, within one day, however the concentration will increase slowly thereafter. Assuming 10 years of continuous operation, N2 concentration would approach 180-540 ppms by the end of the experiment. In the absence of purification, the equivalent of 18-54 ND-GAr volumes of fresh gas would need to be added during the detector live time, to keep the N2 concentration within bounds. Such an estimate represents an upper limit since: i) the outgassing rate of plastics is expected to fall exponentially as the material becomes N2-depleted (eq. 7), and ii) the impact of N2 contamination on the tracking capabilities might be correctable with adequate calibration systems. On the other hand, as discussed earlier, these concentrations are unlikely to create problems for the optical signal.
@@
There are several ways to address the most convenient commissioning strategy, for which a reliable experimental estimate of the N2 outgassing rates would be essential, first and foremost. As an example, the system could be operated in pure argon gas with both hot and cold getters for about 10 days, prior to CF4 injection. N2 concentration would then reach back to 10 ppms in another 10 days for ‘typical’ outgassing values, and 30 days for ‘optimistic’ estimates (eq. 26). Injection of fresh gas in the range 0.3-1% after the first 10 days would guarantee operation around or below the 10 ppm N2 landmark. In the presence of calibration systems, on the other hand, that would allow to correct track distortions as large as 1 cm (100 ppm of N2), few additional ND-GAr volumes, in the range 2-6, would be needed.
@@
As CF4 has a very high GWP, it seems unavoidable that ‘old’ gas needs to be cryorecovered and CF4 distilled prior to disposal to the atmosphere, while fresh gas is being injected.
@@ 5 Summary and conclusions
In this work we have addressed the problem of gas distribution and impurity mitigation for next-generation Time Projection Chambers in the fields of Neutrino Physics and Rare Event Searches. Specifically, we focused on the performance of an Ar/CF4 (99/1) optical TPC of 5 m-length/5 m-diameter at 10 bar, with a field cage covered with Teflon lining, as recently suggested in [14]. We have considered an axial gas distribution scheme (along the electric field direction), that is more independent from the specific end-cap design as well as more easily tractable through computational fluid dynamics (CFD) simulations.
@@
Distributor rods are modeled through a combined approach based on CFD simulations and generalized continuity equations, that allow a straightforward optimization. The approach allows to show that, for the flow rates of relevance to next-generation TPCs, a nearly-perpendicular discharge angle (less than 5 with respect to the normal) can be achieved for hole diameters at the mm-diameter scale, thus ensuring a priori a good bottom-up circulation of the gas for any position along the axial direction, without the creation of gas pockets. Holes are separated by 10-50 mm, the rod diameters being in the range 30-70 mm (possibilities for further reducing this diameter are given in text, although they are outside the scope of this work). The discharge velocity of the holes and the pressure drop at the distributor rod are generally far from the regime where they could pose problems, and are not critical in the optimization.
@@
The proposed gas distribution, based on eight perforated plastic rods (capped at their ends) for injection and eight for extraction, can efficiently sweep the outgassing from the walls when a triple-hole perforation is implemented (one hole situated along the radial direction and two azimuthally). By contrast, single-hole distributors perform worse at sweeping the gas from the walls and are more prone to vortex formation, resulting in much longer times to achieve steady-state conditions. Additionally, the triple-hole configuration leads to an efficient displacement of air during chamber filling (down to 0.1 ppm of O2) in about 42 h time, that is ×$\times 4.5$ faster than expected from a diffusion-dominated 1-D model. (Assuming no pocket formation, the 1-D model provides an estimate of the longest extraction time).
@@
The contamination of O2 (H2O) during chamber operation was studied, assuming circulation in closed loop with the gas forced into conventional ‘cold’ getters. For a realistic flow rate of 6.5 m3/h at 10 bar, a maximum impurity level of 0.3 ppm (3 ppm) would be achieved in a bit over 24 h for typical values of the outgassing rates ($Q_{10h}\simeq 10^{-8}\frac{\textnormal{mbar}\cdot l}{\textnormal{cm}^{2}\cdot\textnormal{s}}$ for O2, and $Q_{10h}\simeq 10^{-7}\frac{\textnormal{mbar}\cdot l}{\textnormal{cm}^{2}\cdot\textnormal{s}}$ for H2O). These concentrations translate into maximum charge losses of about 30% (due to O2) and track distortions of 3 mm (due to H2O). The deterioration in the TPC response is reduced by a factor of 3 after 100 h, approximately following the anticipated reduction of the outgassing rate for technical plastics, as $\sim t^{-1/2}$. Generally speaking, and based on our experimental survey of the Teflon outgassing rates, after 100 h the expected losses due to attachment would be situated in the range 0.5%(‘optimistic’)-10%(‘typical’), with track distortions ranging between 0.3 mm(‘optimistic’) and 1 mm(‘typical’). In the proposed gas distribution scheme, for both O2 and H2O, mixing with the main TPC gas takes place at a pace compatible with the expectation from a diffusion-dominated 1-D model, that provides a benchmark for the shortest mixing time achievable in practice. Importantly, this close-to-ideal mixing renders an uniform distribution of contaminants within merely 24 h, a relevant feature in the need of applying corrections to reconstructed events.
@@
Under the conservative assumption that N2 purification at high flows and in the presence of CF4 needs of dedicated studies and cannot be taken for granted at the moment, we discussed mitigation strategies to the problem of N2 accumulation. It is shown that a total of fresh gas worth 18(‘optimistic’)/54(‘typical’) TPC volumes would be needed during 10 years of continuous operation, a requirement that can be met for instance by injecting fresh gas at a mere 0.3-1% fraction of the flow of the main gas line. If track distortions of up to 1 cm (100 ppm of N2) could be corrected online through (laser-based or other similar) calibration systems, the additional fresh gas will be situated in the range of 2(‘optimistic’)-6(‘typical’) additional TPC volumes. These values are comparable to the ones needed during chamber filling, if TPC pumping would not be possible.
@@
The above discussion, as well as most of the work on the body of the paper, focus on ‘typical’ values of the outgassing and gas flow rates, but there is room for improvement, and lower outgassing values might be achieved through suitable cleaning/surface finish protocols, while higher gas flows will likely become affordable in a few-year time scale. Ultimately this work is aimed at showing that, while operation of the optical TPC discussed here seems perfectly viable from the point of view of material compatibility, detailed outgassing assays of the materials involved in construction seem crucial to ensure the chamber performance and the necessary mitigation strategies.
[…]
@@
This research has been funded by the Spanish Ministry (‘Proyectos de Generación de Conocimiento’, PID2021-125028OB-C21, PID2021-125028OB-C22). It has also received complementary financial support from the regional government Xunta de Galicia (Centro singular de investigación de Galicia accreditation 2019-2022), and by the national program “María de Maeztu” (Units of Excellence program MDM-2016-0692). We thank CERN for computing resources.
@@ Appendix A Iterative implementation of a multi-hole distributor
The mass and momentum -balance equations for each hole section (eqs. 10 and 11) can be solved iteratively: for an assumed gas flow (velocity) at the entrance of the distributor, the pressure upstream of the first hole is scanned till the velocity downstream of the last hole converges to zero (Fig. 16). The set of hole-equations that need to be iteratively propagated along the distributor is:
$\frac{P_{i}+P_{i+1}}{2}$ $=$ $P_{{}_{\text{TPC}}}+\frac{\rho D^{4}(v_{i}-v_{i+1})^{2}}{2d^{4}C_{d,i}^{2}},$ (27)
$P_{i+1}-P_{i}$ $=$ $C_{r,i}\frac{1}{2}\rho(v_{i}^{2}-v_{i+1}^{2}).$ (28)
Cd, Cr are dependent on the gas properties and little-dependent on geometry. We adopted the formalism in [48], in which Cd is parameterized as a function of the ratio $\frac{1}{2}\rho v_{\text{up}}^{2}/\Delta{P}_{\text{hole}}$ and Cr as a function of $\Delta{v}=v_{\text{up}}-v_{\text{down}}$.
Figure 16: Perforated gas distributor capped at its end, together with the main magnitudes introduced in text. The sketch also gives an approximate representation of the solving method, that tries an initial value $P_{\text{up}}$ upstream of the first hole, for an assumed flow, and iterates until a null velocity is obtained at the end of the distributor.
For a complete description, equations 27-28 must be coupled hole-by-hole after accounting for any loss due to friction. The pressure drop in between hole sections is included through Darcy’s friction law:
$P_{\text{up}(i+1)}-P_{\text{down}(i)}=f_{D}\frac{\Delta{L}}{D}\rho\frac{{v_{\text{down}(i)}}^{2}}{2},$ (29)
@@
$Re=v_{\text{down}(i)}\cdot\Delta{L}\cdot\frac{\rho}{\mu},$ (30)

### v2

@@ 4.3 Water contamination
For Teflon, the outgassing rate of water is significantly higher than that of O2, so that if outgassing did not decrease over time then the necessary purity levels might never be achieved. Instead of assuming a constant value, we consider in this case the actual outgassing law based on eq. 2.7, with outgassing rates after 10 h taken from table 2: $3.5\times 10^{-8}$ $\frac{\textnormal{mbar}\cdot{\textnormal{l}}}{\textnormal{cm}^{2}\cdot{\textnormal{s}}}$ (‘optimistic’) and $10\times 10^{-8}$ $\frac{\textnormal{mbar}\cdot{\textnormal{l}}}{\textnormal{cm}^{2}\cdot{\textnormal{s}}}$ (‘typical’).
We concentrate on the case of typical outgassing and flow rates, and five different Teflon thicknesses, including the case where Teflon would be so thick as to preserve a $t^{-1/2}$ behaviour throughout. Based on the analysis in the previous section, the configuration with triple-holes has been chosen. Results are shown in Fig. 13. For thicknesses below 10 mm, a transition between the $t^{-1/2}$ and exponential regimes can be observed. As the outgassing rates of H2O are about 10 times larger than those of O2 in the case of Teflon, the time scales involved are larger too. The horizontal line shows the 1 ppm-level, which is a plausible requirement to keep space resolution undistorted below 1 mm, based on the analysis presented in section 2.1. For a thin-Teflon reflector down to 0.5 mm thickness, coupled to an ESR reflective coating (see e.g. [59]), that landmark value is achieved in 46 h, while it reaches up to 99 h as the Teflon becomes thicker. The result is sobering: when considering the displacement of air during filling and the time needed for water (oxygen) to fall below the concentration that makes its contribution negligible for tracking (attachment), about 6-10 days would be typically needed. Well-defined cleaning protocols and surface treatment, as discussed for example in [41], together with an increase of the gas flow above the levels discussed here, or designing the system in compliance with the requirements needed for pumping, could plausibly reduce these times to the scale of about 1-2 days. Nevertheless, additional (uncontrolled) outgassing sources would render the operation of the detector problematic.
@@
Additional studies are presented in Fig. 14 in order to illustrate the main dependencies. On the left figure, the behaviour for different gases is shown for a fixed flow rate of 6.5 m3/h (argon: continuous lines; xenon: dashed lines). It can be seen how, despite the diffusion coefficient of H2O being larger in argon by about 50% (table 2), the impact on the time-evolution of the impurities is almost indiscernible. When reducing the pressure down to 1 bar, the 10-fold increase in the diffusion coefficient accelerates the mixing as expected; however, the reduction in the position of the maximum is a mere 30%. As the number of normal litres is accordingly reduced by a factor of 10 in those conditions, the concentration of impurities increases in inverse proportion (the Ar and Xe lines at 1 bar have been downscaled by a factor of 10).
Fig. 14-right shows a comparison with the 1-D solutions for the case of time-dependent outgassing (see appendix, section C.2). As for the case of constant outgassing, the results compare well with the 1-D model, although the agreement is qualitatively worse. Artificially increasing the gas diffusion accelerates the mixing, but very mildly in the conditions studied (as indirectly shown in Fig. 14-left). The conclusion is once again that the velocity field in combination with gas diffusion can distribute the outgassed species uniformly over the chamber for the optimised ‘triplet’ distributor, as seen in Fig. 8. Also in this case, simulations for lower (‘optimistic’) outgassing rates display an exact down-scaling compared to the solutions presented in Fig. 14, as expected from the 1-D solutions.
@@ 4.4 Nitrogen contamination
Nitrogen can be removed from the system by flushing the main gas through ‘hot’ (temperature-activated) purifiers/getters, a common asset for the operation of noble gas detectors. The process of nitrogen purification becomes challenging, however, in case of operation under large gas flows (that could modify the getter temperature) and/or in the presence of molecular additives such as CF4 (that may react with the getter materials). Following conventional wisdom, and as long as a purification system compatible with Ar/CF4 operation is not demonstrated at the necessary scale, we shall conservatively assume that N2 accumulates in the TPC as time passes.
@@
Fortunately, the tolerance to N2 is by far the largest of common outgassing species, up to 10 ppm (table 2). Based on the time profile of the outgassing rate (eq. 2.7), the range of outgassing values in table 2, and the chamber volume, the evolution of N2 inside the TPC can be approximated as (see appendix, section C.4):
$f_{\text{N}_{2}}\sim(3-9)\textnormal{ppm}\cdot\sqrt{t[\textnormal{days}]},$ (4.1)
@@
with the pre-factor covering the range dubbed earlier as ‘optimistic’ (3 ppm) and ‘typical’ (9 ppm). In practice, this expression indicates that N2 can approach harmful concentrations quite quickly, within one day; however, the concentration will increase slowly thereafter. Assuming 10 years of continuous operation, N2 concentration would approach 180-540 ppm by the end of the experiment. In the absence of purification, the equivalent of 18-54 ND-GAr volumes of fresh gas would need to be added during the detector lifetime, to keep the N2 concentration within bounds. Such an estimate represents an upper limit since: i) the outgassing rate of plastics is expected to fall exponentially as the material becomes N2-depleted (eq. 2.7), and ii) the impact of N2 contamination on the tracking capabilities might be correctable with adequate calibration systems. Also, as discussed earlier, these concentrations are unlikely to create problems for the optical signal.
@@
There are several ways to address the most convenient commissioning strategy, for which a reliable experimental estimate of the N2 outgassing rates would be essential, first and foremost. As an example, the system could be operated in pure argon gas with both hot and cold getters for about 10 days, prior to CF4 injection. N2 concentration would then reach back to 10 ppm in another 10 days for ‘typical’ outgassing values, and 30 days for ‘optimistic’ estimates (eq. 4.1). Injection of fresh gas in the range of 0.3-1% after the first 10 days would guarantee operation around or below the 10 ppm N2 benchmark. In the presence of calibration systems which would allow for the correction of track distortions as large as 1 cm (100 ppm of N2), few additional ND-GAr volumes (in the range of 2-6) would be needed.
@@
As CF4 has a very high GWP, it seems unavoidable that ‘used’ gas needs to be cryorecovered and CF4 distilled prior to disposal to the atmosphere, whilst fresh gas is being injected.
@@ 5 Summary and conclusions
In this work we have addressed the problem of gas distribution and impurity mitigation for next-generation gas Time Projection Chambers in the fields of Neutrino Physics and Rare Event Searches. Specifically, we focused on the performance of an Ar/CF4 (99/1) optical TPC of 5 m-length/5 m-diameter at 10 bar, with a field cage covered with Teflon lining, as recently suggested in [13]. We have considered an axial gas distribution scheme (along the electric field direction), that is more independent of the specific end-cap design as well as more easily tractable through computational fluid dynamics (CFD) simulations.
@@
Distributor rods are modelled through a combined approach based on CFD simulations and generalised continuity equations, which allow a straightforward optimisation. The approach demonstrates that, for the flow rates of relevance to next-generation TPCs, a nearly-perpendicular discharge angle (less than 5 with respect to the normal) can be achieved for hole diameters at the mm-diameter scale, thus ensuring a priori a good bottom-up circulation of the gas for any position along the axial direction, without the creation of gas pockets. Holes are separated by 10-50 mm, the rod diameters being in the range of 30-70 mm (possibilities for further reducing this diameter are given in the text, although they are outside the scope of this work). The discharge velocity of the holes and the pressure drop at the distributor rod are generally far from the regime where they could pose problems, and are not critical in the optimisation.
@@
The proposed gas distribution, based on eight perforated plastic rods (capped at their ends) for injection and eight for extraction, can efficiently sweep the outgassing from the walls when a triple-hole perforation is implemented (one hole situated along the radial direction and two azimuthally). By contrast, single-hole distributors perform worse at sweeping the gas from the walls and are more prone to vortex formation, resulting in much longer times to achieve steady-state conditions. Additionally, the triple-hole configuration leads to an efficient displacement of air during chamber filling (down to 0.1 ppm of O2) in about 42 h, which is 4.5 times faster than expected from a diffusion-dominated 1-D model. (Assuming no pocket formation, the 1-D model provides an estimate of the longest extraction time).
@@
The contamination of O2 (H2O) during chamber operation was studied, assuming circulation in closed loop with the gas forced into conventional ‘cold’ getters. For a realistic flow rate of 6.5 m3/h at 10 bar, a maximum impurity level of 0.3 ppm (3 ppm) would be achieved in just over 24 h for typical values of the outgassing rates ($Q_{\text{10h}}\simeq 10^{-8}\frac{\textnormal{mbar}\cdot l}{\textnormal{cm}^{2}\cdot\textnormal{s}}$ for O2, and $Q_{\text{10h}}\simeq 10^{-7}\frac{\textnormal{mbar}\cdot l}{\textnormal{cm}^{2}\cdot\textnormal{s}}$ for H2O). These concentrations translate into maximum charge losses of about 30% (due to O2) and track distortions of 3 mm (due to H2O). The deterioration in the TPC response is reduced by a factor of 3 after 100 h, approximately following the anticipated reduction of the outgassing rate for technical plastics, as $\sim t^{-1/2}$. Generally speaking, and based on our experimental survey of the Teflon outgassing rates, after 100 h the expected losses due to attachment would be situated in the range 0.5%(‘optimistic’)-10%(‘typical’), with track distortions ranging between 0.3 mm(‘optimistic’) and 1 mm(‘typical’). In the proposed gas distribution scheme, for both O2 and H2O, mixing with the main TPC gas takes place at a pace compatible with the expectation from a diffusion-dominated 1-D model, which provides a benchmark for the shortest mixing time achievable in practice. Importantly, this close-to-ideal mixing renders a uniform distribution of contaminants within merely 24 h, a relevant feature in the need to apply corrections to reconstructed events.
@@
Under the conservative assumption that N2 purification at high flows and in the presence of CF4 needs dedicated studies and cannot be taken for granted at the moment, we discussed mitigation strategies to the problem of N2 accumulation. It is shown that a total of fresh gas worth 18(‘optimistic’)/54(‘typical’) TPC volumes would be needed during 10 years of continuous operation, a requirement that can be met, for instance, by injecting fresh gas at a mere 0.3-1% fraction of the flow of the main gas line. If track distortions of up to 1 cm (100 ppm of N2) could be corrected online through (laser-based or other similar) calibration systems, the additional fresh gas would be situated in the range of 2(‘optimistic’)-6(‘typical’) additional TPC volumes. These values are comparable to those needed during chamber filling, if TPC evacuation were not possible.
@@
Some considerations towards realism are pertinent. First, a complete 3D simulation of the final system including the end-caps and thermal sources (to account for possible convection currents), as well as the possible thickening of the distributor walls, might suggest variations of the gas distribution proposed, e.g. considering distributor rods of smaller diameter or the possibility of placing them behind the field cage reflector (therefore abandoning the triplet configuration). Second, these conclusions, as well as most of the work in the body of the paper, focuses on ‘typical’ values of the outgassing and gas flow rates, but there is room for improvement, and lower outgassing values might be achieved through suitable cleaning/surface finish protocols, whilst higher gas flows will likely become affordable in a few-year timescale. Ultimately this work is aimed at showing that, while operation of the optical TPC discussed here seems perfectly viable from the point of view of material compatibility, detailed outgassing assays of the materials involved in construction seem crucial to ensure the chamber performance and the necessary mitigation strategies.
Acknowledgments
@@
This research has been funded by the Spanish Ministry (‘Proyectos de Generación de Conocimiento’, PID2021-125028OB-C21, PID2021-125028OB-C22). It has also received complementary financial support from the regional government Xunta de Galicia (Centro singular de investigación de Galicia accreditation 2019-2022), and by the national program “María de Maeztu” (Units of Excellence program MDM-2016-0692). We thank CERN for computing resources. The authors are grateful to the reviewer, whose comments helped improving the manuscript.
@@ Appendix A Iterative implementation of a multi-hole distributor
Results from a three-dimensional CFD simulation performed, illustratively, for a 3-hole distributor, are shown in Fig. 15. The figure provides a cross-section at the holes’ middle plane (the modulus of the gas velocity is represented in the left column, the pressure profile in the right one). Two reference cases are considered: a capped distributor (top) and an open one (bottom). As expected, an open distributor is less efficient, given that only a fraction of the gas exits through the holes and into the chamber—as little as a few % for high gas flow. It also leads to a higher pressure drop along the distributor due to friction, significantly greater directionality of the outgoing gas and a non-uniform discharge pattern. Low efficiency in combination with a directional and non-uniform discharge pattern will adversely affect the distribution of impurities throughout the chamber and the time required to achieve uniform mixing with the main gas. Therefore, we opted in text for a capped distributor, ensuring uniform and normal injection of all the circulated gas.
The mass and momentum -balance equations for each hole section (eqs. 3.1 and 3.2) can be solved iteratively: for an assumed gas flow (velocity) at the entrance of the distributor, the pressure upstream of the first hole is scanned till the velocity downstream of the last hole converges to zero (Fig. 16). The set of hole-equations that need to be iteratively propagated along the distributor is:
$\frac{P_{i}+P_{i+1}}{2}$ $=$ $P_{{}_{\text{TPC}}}+\frac{\rho D^{4}(v_{i}-v_{i+1})^{2}}{2d^{4}C_{d,i}^{2}},$ (A.1)
$P_{i+1}-P_{i}$ $=$ $C_{r,i}\frac{1}{2}\rho(v_{i}^{2}-v_{i+1}^{2}).$ (A.2)
Cd, Cr are dependent on the gas properties and little-dependent on geometry. We adopted the formalism in [52], in which Cd is parameterized as a function of the ratio $\frac{1}{2}\rho v_{\text{u}}^{2}/\Delta{P}_{\text{hole}}$ and Cr as a function of $\Delta{v}=v_{\text{u}}-v_{\text{d}}$.
Figure 16: Perforated gas distributor capped at its end, together with the main magnitudes introduced in text. The sketch also gives an approximate representation of the solving method, that tries an initial value $P_{\text{u}}$ upstream of the first hole, for an assumed flow, and iterates until a null velocity is obtained at the end of the distributor.
For a complete description, equations A.1-A.2 must be coupled hole-by-hole after accounting for any loss due to friction. The pressure drop in between hole sections is included through Darcy’s friction law:
$P_{\text{u}(i+1)}-P_{\text{d}(i)}=f_{D}\frac{\Delta{L}}{D}\rho\frac{{v_{\text{d}(i)}}^{2}}{2},$ (A.3)
@@
$Re=v_{\text{d}(i)}\cdot\Delta{L}\cdot\frac{\rho}{\mu},$ (A.4)